2a-4a=1
What does a equal
-y+6-5y=18
What does y equal
6-z-4z=-4
What does z equal
5(y-6)=-60
What does y equal
-4(x-3)-6x=10
What does x equal
2(z-4)-z=8
What does z equal
-4(n+3)+2n=6

Answer by Nate(3500)

2a-4a=1
-2a=1
a=-1/2
------------------------------------------------------------------------------
-y+6-5y=18
-6y+6=18
-6y=12
y=-2
------------------------------------------------------------------------------
6-z-4z=-4
6-5z=-4
-5z=-10
z=2
------------------------------------------------------------------------------
5(y-6)=-60
5y-30=-60
5y=-30
y=-6
------------------------------------------------------------------------------
-4(x-3)-6x=10
-4x+12-6x=10
-10x=-2
x=1/5
------------------------------------------------------------------------------
2(z-4)-z=8
2z-8-z=8
z=16
------------------------------------------------------------------------------
-4(n+3)+2n=6
-4n-12+2n=6
-2n=18
n=-9